# CNN-Based Vortex Detection in Atomic 2D Bose Gases in the Presence of a Phononic Background

### v1

@@ Abstract
Quantum vortices play a crucial role in both equilibrium and dynamical phenomena in two-dimensional (2D) superfluid systems. Experimental detection of these excitations in 2D ultracold atomic gases typically involves careful labelling of density depletions in absorption images following short time-of-flight expansions, however the presence of a significant phononic background renders the problem challenging, often beyond the capability of simple algorithms or the human eye. Here, we utilize a convolutional neural network (CNN) to detect vortices in the presence of strong long- and intermediate-length scale density modulations in finite-temperature 2D Bose gases. We train the model on datasets obtained from ab initio Monte Carlo simulations using the classical-field method for density and phase fluctuations, and Gross-Pitaevskii simulation of realistic expansion dynamics. We benchmark the performance of our method by comparing it to the matter-wave interferometric detection of vortices, confirming the observed scaling of vortex density across the Berezinskii-Kosterlitz-Thouless (BKT) critical point. The combination of a relevant simulation pipeline with machine-learning methods is a key development towards the comprehensive understanding of complex vortex-phonon dynamics in out-of-equilibrium 2D quantum systems.
@@ 1 Introduction
Quantum vortices are topological excitations found in superfluids and are characterized by integer multiples of ±$\pm 2\pi$ phase winding around a vortex core where the density vanishes. They have been experimentally observed in a wide variety of systems including liquid helium [1, 2], type-II superconductors [3, 4], photon fields [5], exciton-polariton superfluids [6], and in ultracold atom systems with bosonic [7, 8] and fermionic [9] particles. The precise spatial and temporal control of the ultracold atom platform facilitates the study of numerous manifestations of vortex excitations as a consequence of rich static and dynamical many-body phenomena – for example, vortices have been shown to arrange into lattices in rotating Bose-Einstein condensates (BECs) [10] and randomly form following a rapid quench into an ordered phase [11, 12]. An important class of vortex excitations emerges in systems within the 2D XY universality class to which ultracold 2D Bose gases belong [13], having characteristic paired and unpaired configurations across the superfluid transition [14, 15] (see A for a brief overview). The statistics and spatial structure of vortex excitations are one of the direct measures of the many-body states and their time-evolution, hence the precise determination of the number and location of vortices is critical for studying superfluid dynamics. Absorption imaging is typically used to obtain images of the density of ultracold atomic systems [16], however the task of identifying vortices from these is often non-trivial due to their small spatial structure and misleading patterns arising from other excitations in the system. In some situations, it can be simple to count vortices by eye, e.g. in the ground state of 3D rotating BECs where these vortices form an ordered triangular lattice [10, 17, 18]. However, in stationary states of finite-temperature 2D Bose gases the vortices do not form a regular pattern and the presence of phonons, which can result in similar density modulations after time-of-flight expansion [19, 20], makes vortex detection by the human eye very inaccurate and impossible to use for quantitative analysis. Matter-wave interferometry addresses this challenge by taking advantage of the characteristic ±$\pm\pi$ phase discontinuity across vortices (singly quantized) [21, 22]. These distinct features are easy to distinguish from other excitations which have a significantly longer characteristic length scale. This approach has accelerated 2D Bose gas research in recent years by providing a method to obtain the average vortex density from an ensemble of images [22, 23, 24], however this method is limited to measuring averages, missing the rich information available in single-shot detection of quantum vortex locations distributed in the plane of the 2D gas.
@@ 2.1 Numerical simulation of 2D Bose gases
where $\hat{\psi}(\textbf{r})$ and $\hat{\psi}^{\dagger}(\textbf{r})$ are annihilation and creation operators for bosons, respectively, at position vector r in the x-y plane, $g=\widetilde{g}\hbar^{2}/m$ is the quasi-2D interaction parameter, and $V(\textbf{r})$ is the external potential [27]. For the experimental parameters used the trapped gas exhibits near-homogeneous density, therefore, we considered a round box potential of radius 22.5 µm (see B for more experimental details).
@@
We employed the Metropolis-Hastings algorithm [30, 31] to draw random states of a finite-temperature 2D Bose gas from the grand canonical ensemble. We initialized the complex field with zero real and imaginary components. At every iteration, we chose a random site and generated proposed updates to the real and imaginary parts of the on-site field by drawing two random numbers from a normal distribution of zero mean and $\sigma$ standard deviation. We then calculated the grand-canonical weight $p=\exp{\left(-\beta\left[\Delta E-\mu\Delta N\right]/k_{B}T\right)}$, where $\mu$ is the chemical potential, $\Delta E,\Delta N$ are the changes in energy and number of particles in the system by the proposed update of the fields, and $k_{B}$ is the Boltzmann constant. In accordance with the Metropolis-Hastings algorithm, the proposed field update was accepted if $p>1$ or if $p\leq x$, where $x$ was a random number drawn from a $\left[0,1\right]$ uniform distribution.
To ensure thermal equilibrium and uncorrelated fields, we initially ran the algorithm for 200 000 updates per site, and then began sampling fields with a further 5000 updates per site occurring between each field sample. We tuned $\sigma$ such that the mean acceptance rate of field updates was near 0.25, which yields optimal convergence [32, 33]. We collected 2500 fields at eight phase-space densities evenly spaced in the range $\mathcal{D}\in[8.0,11.2]$ by keeping the temperature constant at $T=50$ nK and varying $\mu$. We expected the phase transition to occur at $\mathcal{D}\approx 11$ for our finite-size system which we verified by direct analysis of correlations in C.
@@ 2.2 Data processing
To prepare the output of the Monte Carlo (MC) simulations for training we applied the following steps, laid out in Figure 2, to each field. First, we simulated the short time-of-flight (TOF) expansion, as performed in the experiment, for 5.3 ms which increased the size of the vortex cores. This involved numerically solving the time-dependent Gross-Pitaevskii equation (GPE) [34] using split-step Fourier methods [35]. We performed simulations in 2D and ignored the interactions as they become negligible after the rapid expansion of the cloud in the z-direction which occurs in the experiment. Simulating in 2D captured the important physics while being significantly less computationally demanding than 3D simulations; see D for further discussion.
@@ 3 CNN Design
We designed our CNN with the pixel size of the experimental images (around 1 µm) in mind. Our CNN inputs a 70 $\times$ 70 grayscale image of density, with pixel size 1 µm2, and passes it through four successive convolutional layers, outputting a vortex map, denoted $\widetilde{Y}$, which is a tensor of the same size where the value of each cell is the probability of a vortex being present in the equivalent cell of the input (see Figure 3). The ground-truth, denoted $Y$, is also of the same size, but is restricted to the values 0 and 1.
@@ 4.2 Predictions on experimental images
The experimental dataset of 876 density images, with mean phase-space densities ranging from 8.3 to 16.0, required preprocessing before we could make vortex predictions. First, we removed the random variation of the position of the atom cloud by recentering the images on their center of mass. Then, at each mean phase-space density, we estimated the quasi-condensate radius $r_{\text{QC}}$ by fitting the radial density profile, and defined the region of interest as a circle of radius $0.8r_{\text{QC}}$ as in [22]. This avoided edge effects where the physics is fundamentally different. We found that reasonable predictions required keeping the experimental images unnormalized as a result of the low experimental resolution and two fundamental differences in the appearances of vortices that were not captured by the simulation. Firstly, as phase-space density drops, the fraction of atoms in the non-condensed phase grows which reduces the contrast of density dips. Secondly, the effect of experimental noise relative to signal strength increases as phase-space density decreases. As such, the signal emanating from potential vortex locations must be amplified in order for the method to work. The combination of using unnormalized experimental data and applying image contrast enhancement with morphological operators (see D) achieved this. The models’ predictions are shown in Figure 5.
@@
We could not verify the accuracy of detecting any particular vortex since since their positions are not known a priori in the experimental images. However, we analyzed how the dimensionless vortex number density $\widetilde{n}_{v}=\overline{n}_{v}\xi^{2}$, where $\overline{n}_{v}$ is the mean vortex number density and $\xi$ is the healing length, scales with mean phase-space density $\mathcal{D}_{\text{avg.}}$. The use of $\widetilde{n}_{v}$ accounts for the variation of $\xi$ with $\mathcal{D}_{\text{avg.}}$, hence the characteristic area associated with a vortex. The functional form is expected to follow $\widetilde{n}_{v}\propto\exp({-b\mathcal{D}_{\text{avg.}}})$ with exponential coefficient $b=0.56(5)$ obtained via matter-wave interferometry [22] and supported by simulations. Our fit, shown in Figure 6, yielded $b_{\text{exp}}=0.56(2)$ which is consistent with results using the matter-wave interferometric approach. The presence of the rapid vortex number increase gives us confidence that we have detected vortices, not phonons, since phonon amplitude grows slowly with decreasing $\mathcal{D}_{\text{avg.}}$ and cannot alone explain the phase transition from a superfluid to a normal fluid.
@@ 5 Summary & outlook
We have developed a deep learning-based vortex detection method which can infer vortex locations in density images of 2D Bose gases. The detector is a convolutional neural network (CNN) for which the input is a density image, and the output is a corresponding map of vortex locations. We used Monte Carlo simulations to obtain a synthetic dataset on which to train, circumventing the time-consuming and error-prone task of labeling vortices by hand, and fulfilling the requirement of a large training set for performant machine learning. We have shown that our model performs well when evaluated against out-of-sample simulated images scoring 0.81, 0.65 and 0.72 in precision, recall and F1 metrics, respectively. Further, we have been able to generalize the CNN’s predictive power to experimental images where the predictions are logical and reasonable and reproduce the ground-truth vortex number scaling established in previous work. In this setting, the CNN can distinguish vortices from phonons, and as such it represents a major step toward on-demand, system-wide vortex detection in atomic 2D Bose gases.
The limited resolution of our experimental images necessitated a shallow, size-preserving CNN. With better image quality, the vortices will be more readily detectable, and at the same time larger CNNs involving maxpooling layers can be leveraged to significantly boost confidence in predictions. Then, applications to real-time tracking of free vortices as described in [39, 40, 41], or vortex pairing dynamics as in [36] can be considered. Also, we hope such improvements will allow this approach to extend to bilayer systems such as those in [22, 24, 20], where inter-layer interactions are expected to give rise to correlated free vortices in the two layers, which can be distinguished from other vortices by density depletions of different magnitude.
@@ Appendix B Experimental details
Appendix C Simulation verification
We confirmed that our simulation accurately captured both the superfluid and normal phases and the transition between them, by analyzing the two-point phase correlation function. We can see from Figure 8(a) that the BKT transition occurs around $\mathcal{D}\approx 11$ as evidenced by the narrow crossover region between the root-mean-square errors of algebraic and exponential fits to the correlation function. In Figure 8(b), we see that the crossover region places the algebraic exponent $\eta$, defined when fitting $r^{-\eta}$ to the correlation function, just below the universal value of $0.25$ found in the thermodynamic limit – a typical result for finite-sized systems [22].
@@ Appendix E Neural network architecture and training details
Our CNN has four successive convolutional layers of which the first three use the rectified linear unit (ReLU) activation function and the final layer uses sigmoid to obtain a probability between zero and one. Each convolutional layer, barring the final one, is followed by a batch normalization layer to stabilize and speed up training [51]. The filter sizes and number of filters in each layer are shown in Table LABEL:table:nn_architecture. Including the batch normalization layers, there are 24 406 learnable parameters.

### v2

@@ Abstract
Quantum vortices play a crucial role in both equilibrium and dynamical phenomena in two-dimensional (2D) superfluid systems. Experimental detection of these excitations in 2D ultracold atomic gases typically involves examining density depletions in absorption images, however the presence of a significant phononic background renders the problem challenging, beyond the capability of simple algorithms or the human eye. Here, we utilize a convolutional neural network (CNN) to detect vortices in the presence of strong long- and intermediate-length scale density modulations in finite-temperature 2D Bose gases. We train the model on datasets obtained from ab initio Monte Carlo simulations using the classical-field method for density and phase fluctuations, and Gross-Pitaevskii simulation of realistic expansion dynamics. We use the model to analyze experimental images and benchmark its performance by comparing the results to the matter-wave interferometric detection of vortices, confirming the observed scaling of vortex density across the Berezinskii-Kosterlitz-Thouless (BKT) critical point. The combination of a relevant simulation pipeline with machine-learning methods is a key development towards the comprehensive understanding of complex vortex-phonon dynamics in out-of-equilibrium 2D quantum systems.
@@ 1 Introduction
Quantum vortices are topological excitations found in superfluids and are characterized by integer multiples of ±$\pm 2\pi$ phase winding around a vortex core where the density vanishes. They have been experimentally observed in a wide variety of systems including liquid helium [1, 2], type-II superconductors [3, 4, 5, 6], photon fields [7], exciton-polariton superfluids [8, 9], and in ultracold atom systems with bosonic [10, 11, 12] and fermionic [13, 14, 15] particles. The precise spatial and temporal control of the ultracold atom platform facilitates the study of numerous manifestations of vortex excitations as a consequence of rich static and dynamical many-body phenomena – for example, vortices have been shown to arrange into lattices in rotating Bose-Einstein condensates (BECs) [16] and randomly form following a rapid quench into an ordered phase [14, 15, 17, 18, 19]. An important class of vortex excitations emerges in systems within the 2D XY universality class to which ultracold 2D Bose gases belong [20], having characteristic paired and unpaired configurations across the superfluid transition [21, 22] (see A for a brief overview). The statistics and spatial structure of vortex excitations are one of the direct measures of the many-body states and their time-evolution, hence the precise determination of the number and location of vortices is critical for studying superfluid dynamics. Absorption imaging is typically used to obtain images of the density of ultracold atomic systems [23, 24], however the task of identifying vortices from these is often non-trivial due to their small spatial structure and misleading patterns arising from other excitations in the system. In some situations, it can be simple to count vortices by eye, e.g. in the ground state of 3D rotating BECs where these vortices form an ordered triangular lattice [16, 25, 26]. However, in stationary states of finite-temperature 2D Bose gases the vortices do not form a regular pattern and the presence of phonons, which can result in similar density modulations after time-of-flight expansion [27, 28], makes vortex detection by the human eye very inaccurate and impossible to use for quantitative analysis. Matter-wave interferometry addresses this challenge by taking advantage of the characteristic ±$\pm\pi$ phase discontinuity across vortices (singly quantized) [11, 12, 29]. These distinct features are easy to distinguish from other excitations which have a significantly longer characteristic length scale. This approach has accelerated 2D Bose gas research in recent years by providing a method to obtain the average vortex density from an ensemble of images [29, 30, 31], however this method is limited to measuring averages, missing the rich information available in single-shot detection of quantum vortex locations distributed in the plane of the 2D gas.
@@ 2.1 Numerical simulation of 2D Bose gases
where $\hat{\psi}(\textbf{r})$ and $\hat{\psi}^{\dagger}(\textbf{r})$ are annihilation and creation operators for bosons, respectively, at position vector r in the x-y plane, $g=\widetilde{g}\hbar^{2}/m$ is the quasi-2D interaction parameter, $\widetilde{g}$ is the dimensionless 2D interaction strength, and $V(\textbf{r})$ is the external potential [34]. For the experimental parameters used the trapped gas exhibits near-homogeneous density, therefore, we considered a round box potential of radius 22.5 µm. This also allowed us to define the phase-space density of the gas as $\mathcal{D}=nh^{2}/(2\pi mk_{B}T)\propto n/T$, where $n$ is the average 2D density, and $T$ is the temperature (see B for more experimental details).
@@
We employed the Metropolis-Hastings algorithm [37, 38] to draw random states of a finite-temperature 2D Bose gas from the grand canonical ensemble. We initialized the complex field with zero real and imaginary components. At every iteration, we chose a random site and generated proposed updates to the real and imaginary parts of the on-site field by drawing two random numbers from a normal distribution of zero mean and $\sigma$ standard deviation. We then calculated the grand-canonical weight $p=\exp{\left(-\beta\left[\Delta E-\mu\Delta N\right]/k_{B}T\right)}$, where $\mu$ is the chemical potential, $\Delta E,\Delta N$ are the changes in energy and number of particles in the system by the proposed update of the fields, and $k_{B}$ is the Boltzmann constant. In accordance with the Metropolis-Hastings algorithm, the proposed field update was accepted if $p>1$ or if $p\leq x$, where $x$ was a random number drawn from a $\left[0,1\right]$ uniform distribution. Full details can be found in C.
@@ 2.2 Data processing
To prepare the output of the Monte Carlo (MC) simulations for training we applied the following steps, laid out in Figure 2, to each field. First, we simulated the short time-of-flight (TOF) expansion, as performed in the experiment, for 5.3 ms which increased the size of the vortex cores. This involved numerically solving the time-dependent Gross-Pitaevskii equation (GPE) [39, 40] using split-step Fourier methods [41]. We performed simulations in 2D and ignored the interactions as they become negligible after the rapid expansion of the cloud in the z-direction which occurs in the experiment. Simulating in 2D captured the important physics while being significantly less computationally demanding than 3D simulations; see D for further discussion.
@@ 3 CNN Design
We designed our CNN with the pixel size of the experimental images (around 1 µm) in mind. Our CNN inputs a 70 $\times$ 70 grayscale image of density, with pixel size 1 µm2, and passes it through four successive convolutional layers, outputting a vortex map, denoted $\widetilde{Y}$, which is a tensor of the same size where the value of each cell is the probability of a vortex being present in the equivalent cell of the input (see Figure 3). The ground-truth, denoted $Y$, is also of the same size, but is restricted to the values 0 and 1. The number of convolutional layers and filters was iteratively optimized to capture vortex patterns with high accuracy while mitigating overfitting to the training data. This refinement ensured robust generalization to the experimental data while minimizing the computational overhead associated with over-parameterization. Notably, omitting fully connected layers significantly reduced memory requirements, improving the model’s efficiency and adaptability for deployment.
@@ 4.2 Predictions on experimental images
The experimental dataset had 876 density images, with mean phase-space densities ranging from 8.3 to 16.0, and required preprocessing before we could make vortex predictions. First, we removed the random variation of the position of the atom cloud by recentering the images on their center of mass. Then, at each mean phase-space density, we estimated the quasi-condensate radius $r_{\text{QC}}$ by fitting the radial density profile, and defined the region of interest as a circle of radius $0.8r_{\text{QC}}$ as in [29]. This avoided edge effects where the physics is fundamentally different. We found that reasonable predictions required keeping the experimental images unnormalized as a result of the low experimental resolution and two fundamental differences in the appearances of vortices that were not captured by the simulation. Firstly, as phase-space density drops, the fraction of atoms in the non-condensed phase grows which reduces the contrast of density dips. Secondly, at lower phase-space densities, the reduced detected atom number leads to a lower signal-to-noise ratio. As such, the signal emanating from potential vortex locations must be amplified in order for the method to work. The combination of using unnormalized experimental data and applying image contrast enhancement with morphological operators (see D) achieved this. The models’ predictions are shown in Figure 5.
@@
We could not verify the accuracy of detecting any particular vortex since their positions are not known a priori in the experimental images. However, we analyzed how the dimensionless vortex number density $\widetilde{n}_{v}=\overline{n}_{v}\xi^{2}$, where $\overline{n}_{v}$ is the mean vortex number density and $\xi$ is the healing length, scales with mean phase-space density $\mathcal{D}_{\text{avg.}}$. The use of $\widetilde{n}_{v}$ accounts for the variation of $\xi$ with $\mathcal{D}_{\text{avg.}}$, hence the characteristic area associated with a vortex. The functional form is expected to follow $\widetilde{n}_{v}\propto\exp({-b\mathcal{D}_{\text{avg.}}})$ with exponential coefficient $b=0.56(5)$ obtained via matter-wave interferometry [29] and supported by simulations. Our fit, shown in Figure 6, yielded $b_{\text{exp}}=0.56(2)$ which is consistent with results using the matter-wave interferometric approach. The presence of the rapid vortex number increase gives us confidence that we have detected vortices, not phonons, since phonon amplitude grows slowly with decreasing $\mathcal{D}_{\text{avg.}}$ and cannot alone explain the phase transition from a superfluid to a normal fluid.
@@ 5 Summary & outlook
We have developed a deep learning-based vortex detection method which can infer vortex locations in density images of 2D Bose gases. The detector is a convolutional neural network (CNN) for which the input is a density image, and the output is a corresponding map of vortex locations. We used Monte Carlo simulations to obtain a synthetic dataset on which to train, circumventing the time-consuming and error-prone task of labeling vortices by hand, and fulfilling the requirement of a large training set for performant machine learning. We have shown that our model performs well when evaluated against out-of-sample simulated images scoring 0.81, 0.65 and 0.72 in precision, recall and F1 metrics, respectively. Further, we have been able to generalize the CNN’s predictive power to experimental images where vortices cannot be identified by simple algorithms or the human eye. These predictions were validated against a standard matter-wave interferometric benchmark from previous work, successfully reproducing the ground-truth vortex number scaling. Our results demonstrate that the CNN reliably distinguishes vortices from phonons, marking a significant step toward on-demand, system-wide vortex detection in atomic 2D Bose gases.
With improved image quality, the vortices will be more readily detectable, allowing simple and efficient models—such as the one presented in this paper—to make predictions with even greater confidence and accuracy. Furthermore, transformers can be used to enhance existing CNN-based approaches by integrating global context, better capturing the relationships between vortices, though this comes with increased computational demands. Looking ahead, applications to real-time tracking of free vortices as described in [45, 46, 47], or vortex pairing dynamics as in [42] will become more feasible. We hope such improvements will allow this approach to extend to bilayer systems such as those in [29, 31, 28], where inter-layer interactions are expected to give rise to correlated free vortices in the two layers, which can be distinguished from other vortices by density depletions of different magnitude.
@@ Appendix B Experimental details
Appendix C Simulation details & verification
To ensure thermal equilibrium and uncorrelated fields, we initially ran the algorithm for 200 000 updates per site, and then began sampling fields with a further 5000 updates per site occurring between each field sample. We tuned $\sigma$ such that the mean acceptance rate of field updates was near 0.25, which yields optimal convergence [56, 57]. We collected 2500 fields at eight phase-space densities evenly spaced in the range $\mathcal{D}\in[8.0,11.2]$ by keeping the temperature constant at $T=50$ nK and varying $\mu$. We expected the phase transition to occur at $\mathcal{D}\approx 11$ for our finite-size system which we verify by direct analysis of the two-point phase correlation function in Figure 8
We can see from Figure 8(a) that the BKT transition occurs around $\mathcal{D}\approx 11$ as evidenced by the narrow crossover region between the root-mean-square errors of algebraic and exponential fits to the correlation function. In Figure 8(b), we see that the crossover region places the algebraic exponent $\eta$, defined when fitting $r^{-\eta}$ to the correlation function, just below the universal value of $0.25$ found in the thermodynamic limit – a typical result for finite-sized systems [29].
@@ Appendix E Neural network architecture and training details
Our CNN has four successive convolutional layers of which the first three use the rectified linear unit (ReLU) activation function and the final layer uses sigmoid to obtain a probability between zero and one. Each convolutional layer, barring the final one, is followed by a batch normalization layer to stabilize and speed up training [59]. The filter sizes and number of filters in each layer are shown in Table 1. Including the batch normalization layers, there are 24 406 learnable parameters.